OpenStudy (anonymous):

Adele had a 5/6 of a foot long board and used 1/2 of a foot for her shop project. What length of board did she have left?

OpenStudy (anonymous):

\[\frac{ 5 }{ 6 }-\frac{ 1 }{ 2 }=\frac{ 5 }{ 6 }-\frac{ 1 }{ 2 }\times \frac{ 3 }{ 3 }=\frac{ 5 }{ 6 }-\frac{ 3 }{ 6 }=\frac{ 2 }{ 6 }=\frac{ 1 }{ 3 }\]